Relationships between Categorical Variables

Top of Form


A study of 451 randomly selected adults asked "Are you in favor or opposed to the death penalty for persons convicted of murder?" Show your fraction and the percent (rounded to one decimal place).

White

Black

Hispanic

total

Favor

212

55

56

323

Oppose

56

29

43

128

total

268

84

99

451


A. What percent of all respondents are in favor of the death penalty?

B. What percent of all respondents are white adults that are opposed to the death penalty?

C. What is the percent marginal distribution of race?

D. What is the percent marginal distribution of death penalty stance?

E. What percent of Hispanics oppose the death penalty?

F. What percent of those opposing the death penalty are Hispanic?

        G. What does the percentage 56 / 268 = 20.9% represent? Interpret this in one sentence.

H. What is the conditional distribution of death penalty stance by race?

I. Are race and position on death penalty independent? Explain your response using the appropriate percentages. (Your response must include the right percentages and interpretations)
Answer #1
white balck hispanic total
favor 212 55 56 323
oppose 56 29 43 128
total 268 84 99 451

a) percent of all respondents are in favor of the death penalty = 323/451*100=                           71.62%
                          
b) percent of all respondents are white adults that are opposed to the death penalty=56/451=                           12.42%
                          
c) marginal distribution of white = 268/451=                           59.42%
                          
marginal distribution of Black = 84/451=                           18.63%
                          
marginal distribution of Hispanci=99/451=                           21.95%
                          
d) marginal distribution of favor = 323/451=                           71.62%
                          
marginal distribution of who oppose=128/451=                           28.38%
--------------------------------------------------------------

please note that only first four subparts of a ques per post can be answered as per HOMEWORKLIB RULES